OpenStudy (anonymous):

The measure of angle B is 65.8 degrees. Find the measure of the complement and supplement of angle B.

OpenStudy (anonymous):

complement = 90 - 65.8 = 24.2 degrees, supplement = 180 - 65.8 = 114.2 degrees

OpenStudy (cggurumanjunath):